A committee contains 8 members to be chossen from 9 teachers and 4 students.
How many ways can this be done if the committee is to contain exactly 3
students?

Found 3 solutions by TimothyLamb, stanbon, Edwin McCravy:
Answer by TimothyLamb(4379)

the committee has 5 teachers and 3 students:
---
9*8*7*6*5 * 4*3*2 = 362880
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by stanbon(75887)

a comeetie contains 8members is to be chossen for 9teachers and 4students.How many ways can be done if comeetie contains exactly 3students.
------------------
# of ways to succeed: 4C2*9C6 = 6*84 - 504
------------------
# of possible groups of 8:: 13C8 = 1287
-----
P(condition) = 504/1287 = 0.3916
----------------
Cheers,
Stan H.
----------------

Answer by Edwin McCravy(20054)

Each committee of 8 contains exactly 3 students and 5 teachers.

Choose the 3 students C(4,3) = 4 ways.
For each of those 4 ways, we can choose the 5 teachers C(9,5) = 126 ways.

Answer: 4×126 = 504 ways.

Edwin